Solve each equation for y.
y + 1/2 = -1/3 (x + 1/2)

ISBN: 13-978-0-07-326862-0

Answer by checkley71(8403)

Y+1/2=-1/3(X+1/2) OR Y=-1/2-(1/3)X-1/6 OR Y=-2/3-(1/3)X